OpenStudy (anonymous):

-4x + 5(x - 2)= -16 + 10 what does x equal

OpenStudy (constanceleak):

-6=x

OpenStudy (constanceleak):

nevermind

OpenStudy (anonymous):

-4x+5(x-2) = -16+10 -4x+5(x-2) = -6 -4x+[5x+5(-2)] = -6 -4x+5x+(-10) = -6 -4x+5x = -4 1x = -4 x = -4